# Fitted value shrinkage

### v5

@@ 1 Introduction
Fitting (1) by penalized least-squares has been studied by many scholars. Well-studied penalties include the ridge penalty (Hoerl and Kennard, 1970), the bridge/lasso penalty (Frank and Friedman, 1993; Tibshirani, 1996), the adaptive lasso penalty (Zou, 2006), the SCAD penalty (Fan and Li, 2001), and the MCP penalty (Zhang, 2010). Unfortunately, these methods’ fitted values are not invariant to invertible linear transformations of $X$. This is particularly problematic when categorical variables (with three or more categories) and their interactions are encoded in $X$ because a change in the coding can change the fit. The Group Lasso (Yuan and Lin, 2006) and its variants with different standardization (Choi et al., 2012; Simon and Tibshirani, 2012) could be used when there are categorical predictors, but these methods still lack invariance and our simulation and data examples illustrate their instability. This lack of invariance is also present in principal components regression (Hotelling, 1933), and partial least squares (Wold, 1966).
Invariance to invertible linear transformations is important in any application that includes categorical predictors and their interactions: the model fit should not change by changing the coding scheme used to represent these variables in the design matrix. We illustrate that a change in the reference level coding of categorical predictors (see Section 6.2) and that a rotational transformation of the design matrix that removes zeros from the regression coefficient vector (see Section 6.3 and 6.5) leads the regression shrinkage methods that lack this invariance to perform worse. In contrast, our method performed the same before and after these transformations.
@@ 2.1 Method description
We will derive an optimal value of $\gamma$ that minimizes $\mathbb{E}\|X\hat{\beta}^{(\gamma)}-X\beta\|^{2}$ and propose two estimators of it: one for low dimensions and one for high dimensions. We also explore using cross validation to select $\gamma$ when the response and predictor measurement pairs are drawn from a joint distribution. Conveniently, our results generalize to shrinkage towards a submodel’s fitted values $P_{X_{0}}Y$, where $X_{0}$ is a matrix with a proper subset of the columns of $X$.
@@ 4.3 Tuning parameter selection in high dimensions
5 Shrinking toward the fitted values of a submodel
The previous sections developed our new shrinkage method for linear regression with fitted values that are invariant to invertible linear transformations of the design matrix $X$. Its shrinkage target was the fitted values of the intercept-only model $P_{1}Y$. We can generalize this so that the shrinkage target is $P_{X_{0}}Y$, where $X_{0}$ is the design matrix for a submodel formed from a proper subset of the columns of $X$, e.g. $X_{0}=1_{n}$ (the first column of $X$) as it was previously. We consider the following generalization of (1):
$X\tilde{\beta}^{(\gamma)}=\gamma P_{X}Y+(1-\gamma)P_{X_{0}}Y.$
Proposition 6.
For all $(n,p)\in\{1,2,\ldots\}\times\{1,2,\ldots\}$,
We can estimate $\tilde{\gamma}_{\rm opt}$ in low dimensions with
$\tilde{\gamma}=(1-1/\tilde{F})\cdot 1(\tilde{F}>1),$ (19)
Recall that $r={\rm rank}(X)$ and let $r_{0}={\rm rank}(X_{0})$.
Proposition 7.
Proposition 8.
@@ 7.2 Low dimensional data analyses with categorical variables and their interactions
We analyzed two data sets from existing R packages to illustrate the performance of our estimators when categorical variables with interactions are present in the model. The competitors and setup for the data experiments are nearly identical to the previous Section 6.2, except we added three new fitted-value shrinkage estimators that shrink toward the submodel without interactions instead of the intercept-only model. These new submodel shrinkage methods are labeled 2n-Repsb, 2n-Gsb, 2n-Essb, 2n-Es95sb, and they respectively correspond to 2n-Rep, 2n-G, 2n-Es, and 2n-Es95.
@@
Table 2: The performance comparison table for two data sets in Section 7.2 with two different coding strategies. The values are mean squared prediction errors averaged over 50 training/test splits. The numbers in parentheses are normalized sample standard deviations. The labels are defined in Section 6.1 except that last three are illustrated in Section 7.2. The suffix ”sb” for each estimator corresponds to shrinking towards the submodel without interactions instead of shrinking towards the intercept-only model (see Section 5 for details). Boldface indicates the best model. Underlined are our main estimator and its submodel shrinkage variant.
@@ 7.3 High dimensional data experiments
For high-dimensional data examples, we randomly selected subjects from existing data sets so that there were fewer subjects than predictors. We used the same splitting and evaluation procedure that we used in Sections 7.1 and 7.2. The competitors are same as those considered in Section 6.4 except that we excluded 2n-Rep4 which had nearly identical performance to 2n-Rep3 in the simulation study. The following is a description of the examples:
@@
8 Future directions
@@ Acknowledgement
.7 Additional simulations: submodel shrinkage
We conduct additional simulations which is again the linear regression subjects model (1) with $n=100,\,p=75$. The design matrix $X$ has ones in its first column and independent draws from $N_{p-1}(0,\Sigma)$ in the remaining entries on each row, where $\Sigma_{jk}=0.5^{|j-k|}$. However, in this section we explore, possible scenarios where the shrinkage towards known submodel (or submatrix) $X_{0}$ has an advantage over the shrinkage towards the intercept only model (see Section 5 for details). Let $X_{0}$ denote the sub-matrix of $X$ which consists of the first $p_{0}$ columns of $X$. We denote $X=(X_{0},X_{1})$. We consider $p_{0}\in\{5,10,15,20,25\}$. And, we update $X_{1}$ by $X_{1}^{*}=Q_{X_{0}}X_{1}=(I-P_{X_{0}})X_{1}$, and further replace $(X_{0},X_{1})$ with $(X_{0},X_{1}^{*})$ for $X$. We conduct this orthogonalization procedure to decompose terms easier (see (41) below). Then we create

### v6

@@ 1 Introduction
Fitting (1) by penalized least-squares has been studied by many scholars. Well-studied penalties include the ridge penalty (Hoerl and Kennard, 1970), the bridge/lasso penalty (Frank and Friedman, 1993; Tibshirani, 1996), the adaptive lasso penalty (Zou, 2006), the SCAD penalty (Fan and Li, 2001), and the MCP penalty (Zhang, 2010). Unfortunately, these methods’ fitted values are not invariant to invertible linear transformations of $X$. The lack of invariance is particularly problematic when a practitioner wants to find a set of linear combinations of a subset of the predictors to reduce multicollinearity, because a rotational change in the model matrix can substantially change the predicted values of the penalized methods that lack invariance. In addition, invariance to invertible linear transformations is also important in any application that includes categorical predictors (with three or more categories) and their interactions: the model fit should not change by changing the coding scheme used to represent these variables in the design matrix. The Group Lasso (Yuan and Lin, 2006) and its variants with different standardization (Choi et al., 2012; Simon and Tibshirani, 2012) could also be used when there are categorical predictors, but these methods still lack invariance and our simulation and data examples illustrate their instability. This lack of invariance is also present in principal components regression (Hotelling, 1933), and partial least squares (Wold, 1966).
We illustrate that a change in the reference level coding of categorical predictors (see Section 5.2) and that a rotational transformation of the design matrix that removes zeros from the regression coefficient vector (see Section 5.3 and 5.5) leads the regression shrinkage methods that lack this invariance to perform worse. In contrast, our method performed the same before and after these transformations. We also propose a method to estimate the regression’s error variance in high dimensions that may be of independent interest.
@@ 2.1 Method description
We will derive an optimal value of $\gamma$ that minimizes $\mathbb{E}\|X\hat{\beta}^{(\gamma)}-X\beta\|^{2}$ and propose two estimators of it: one for low dimensions and one for high dimensions. We also explore using cross validation to select $\gamma$ when the response and predictor measurement pairs are drawn from a joint distribution. Conveniently, our results generalize to shrinkage towards a submodel’s fitted values $P_{X_{0}}Y$, where $X_{0}$ is a matrix with a proper subset of the columns of $X$ (see Section Shrinking toward the fitted values of a submodel of the Supplementary material).
@@ 5.4 High-dimensional experiments
2n-Rep1, 2n-Rep3, and 2n-Rep4 are based on our variance estimator (15); and 2n-Rep2 is based on the variance estimator proposed by Liu et al. (2020). From Proposition 4, 2n-Rep1 is suitable when $\delta^{2}=\|\mu-P_{1}\mu\|^{2}$ is large relative to $n\sigma^{2}$, e.g. $\gamma_{\rm opt}\rightarrow 1$. In contrast, based on Proposition 4 and 5, 2n-Rep2, 2n-Rep3, and 2n-Rep4 are suitable when $\delta^{2}=\|\mu-P_{1}\mu\|^{2}$ is small relative to $n\sigma^{2}$, e.g. $\gamma_{\rm opt}\rightarrow 0$. As we will illustrate below, 2n-Rep3 generally outperforms 2n-Rep2 and 2n-Rep4. So we recommend using 2n-Rep1 for stronger signals and 2n-Rep3 for weaker signals.
@@ 6.2 Low dimensional data analyses with categorical variables and their interactions
We analyzed two data sets from existing R packages to illustrate the performance of our estimators when categorical variables with interactions are present in the model. The competitors and setup for the data experiments are nearly identical to the previous Section 5.2, except we added three new fitted-value shrinkage estimators that shrink toward the submodel without interactions instead of the intercept-only model. We refer the readers to Section Shrinking toward the fitted values of a submodel of the Supplementary material for the definition of the submodel shrinkage. These new submodel shrinkage methods are labeled 2n-Repsb, 2n-Gsb, 2n-Essb, 2n-Es95sb, and they respectively correspond to 2n-Rep, 2n-G, 2n-Es, and 2n-Es95.
@@
Table 2: The performance comparison table for two data sets in Section 6.2 with two different coding strategies. The values are mean squared prediction errors averaged over 50 training/test splits. The numbers in parentheses are normalized sample standard deviations. The labels are defined in Section 5.1 except that last three are illustrated in Section 6.2. The suffix ”sb” for each estimator corresponds to shrinking towards the submodel without interactions instead of shrinking towards the intercept-only model (see Section Shrinking toward the fitted values of a submodel for details). Boldface indicates the best model. Underlined are our main estimator and its submodel shrinkage variant.
@@ 6.3 High dimensional data experiments
For high-dimensional data examples, we randomly selected subjects from existing data sets so that there were fewer subjects than predictors. We used the same splitting and evaluation procedure that we used in Sections 6.1 and 6.2. The competitors are same as those considered in Section 5.4 except that we excluded 2n-Rep4 which had nearly identical performance to 2n-Rep3 in the simulation study. The following is a description of the examples:
@@
7 Discussion
Lasso and ridge-penalized least squares are popular and powerful in practice. However, their fitted values lack invariance to invertible linear transformations of the design matrix, which is undesirable when there are categorical predictors and interactions. Our simulation studies and data analyses illustrated that our proposed method performed comparably to ridge-penalized least squares, and so we recommend that practitioners use our method in any problem that they would use ridge-penalized least squares. Our method serves as a companion to ridge-penalized least squares, with the advantage of preserving invariance to invertible linear transformations of the design matrix.
@@ Acknowledgement
Shrinking toward the fitted values of a submodel
In Section 24, we developed our new shrinkage method for linear regression with fitted values that are invariant to invertible linear transformations of the design matrix $X$. Its shrinkage target was the fitted values of the intercept-only model $P_{1}Y$. We can generalize this so that the shrinkage target is $P_{X_{0}}Y$, where $X_{0}$ is the design matrix for a submodel formed from a proper subset of the columns of $X$, e.g. $X_{0}=1_{n}$ (the first column of $X$) as it was previously. We consider the following generalization of (1):
$X\tilde{\beta}^{(\gamma)}=\gamma P_{X}Y+(1-\gamma)P_{X_{0}}Y.$
Proposition 6.
For all $(n,p)\in\{1,2,\ldots\}\times\{1,2,\ldots\}$,
We can estimate $\tilde{\gamma}_{\rm opt}$ in low dimensions with
$\tilde{\gamma}=(1-1/\tilde{F})\cdot 1(\tilde{F}>1),$ (38)
Recall that $r={\rm rank}(X)$ and let $r_{0}={\rm rank}(X_{0})$.
Proposition 7.
Proposition 8.
.1 Additional simulations: submodel shrinkage
We conduct additional simulations which is again the linear regression subjects model (1) with $n=100,\,p=75$. The design matrix $X$ has ones in its first column and independent draws from $N_{p-1}(0,\Sigma)$ in the remaining entries on each row, where $\Sigma_{jk}=0.5^{|j-k|}$. However, in this section we explore, possible scenarios where the shrinkage towards known submodel (or submatrix) $X_{0}$ has an advantage over the shrinkage towards the intercept only model (see Section Shrinking toward the fitted values of a submodel for details). Let $X_{0}$ denote the sub-matrix of $X$ which consists of the first $p_{0}$ columns of $X$. We denote $X=(X_{0},X_{1})$. We consider $p_{0}\in\{5,10,15,20,25\}$. And, we update $X_{1}$ by $X_{1}^{*}=Q_{X_{0}}X_{1}=(I-P_{X_{0}})X_{1}$, and further replace $(X_{0},X_{1})$ with $(X_{0},X_{1}^{*})$ for $X$. We conduct this orthogonalization procedure to decompose terms easier (see (41) below). Then we create